A particle of mass 1kg is moving along a straight line y=x+4. Both x and y are in metres. Velocity of the particle is 2m/s. Find the magnitude of angular momentum of the particle about the origin

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Magnitude of angular momentum= mvr

m=1 kg

v=2 m/s

r=4 m

so L=8 kg m2/s

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com



vikas askiitian expert
13 years ago

y = x +4

x-y+4 =0

perpendicular distance of this line from origin is given by

(x-y+4)/21/2 =4/21/2  = d

angulat momentam = mvd

                            =8/21/2 =4sqrt2 units